OpenStudy (anonymous):

simplify 3-x/3x/9-x^2/5x^3

OpenStudy (campbell_st):

so you are dividing by a fraction... so find the reciprocal of the denominator and multiply \[\frac{3 - x}{3x} \times \frac{5x^3}{9 - x^2}\] you can factor the denominator of the 2nd fraction so \[\frac{3 - x}{3x} \times \frac{5x^3}{(3 - x)(3 + x)}\] this will help you to simplify the expression.

OpenStudy (campbell_st):

ummm removing the obvious common factor leaves \[\frac{5x^3}{3x(3 + x)}\] there is still a common factor that can be removed...

OpenStudy (campbell_st):

yes there is a common factor of x that can be removed... just distribute the denominator for the final solution.